3. If 50 of the 500,000 people in Atlanta have PKU, how many people are carriers?

4. In Atlanta, if a woman with PKU marries a man without symptoms of PKU, what is the chance that they will have a child with PKU?
Answer #1

Answer:

3).

Frequency of PKU trait = q^2 = 50/500000 = 0.0001

Frequency of PKU allele = q = SQRT of 0.0001 = 0.01

Frequency of normal allele = p = 1-0.01 = 0.99 (as p+q=1)

Frequency of carriers = 2pq= 2*0.99*0.01 = 0.0198

Number of carriers = 0.0198 * 500000 = 9900

4).

PKU = p

Normal = P

Based on the information, the man would be the carrier. Then only the child would have the PKU trait.

pp (woman) x (man) Pp ---Parents

P

p

p

Pp (normal)

pp (PKU)

The chance of child with PKU = 50%